Can you please help me with this work problem.
Thank you!!
A rectangular observation deck overlooking a scenic valley is to be built. It is to have dimensions 6 meters by 12 meters. For commercial reasons it is decided to add on a rectangular shelter which would take up 40 square meters of the deck. The remainder of the deck will be used as a walkway. It will have uniform width. How wide should this walkway be?

Answer by richwmiller(17219)

I just solved this.
6*12=72
72-40=32
the shelter will be 10by 4
the walkway will be 2 meters wide.
and take up 32 m^2
10by 2 for one side of the shelter and 4*2 for the other side and 2*2 for the corner
(10+2+4)*2=32